OpenStudy (anonymous):

Differential Equations Find the largest interval which includes x=0 for which the given initial-value problem has a unique solution. (Enter your answer in interval notation) (x-3)y"+2y=x, y(0)=0, y'(0)=1

OpenStudy (anonymous):

what are your choices

OpenStudy (anonymous):

...

OpenStudy (anonymous):

invalid question @mathmale

OpenStudy (mathstudent55):

\((-\infty, \infty\ )\)

OpenStudy (agent0smith):

Intriguing question.

OpenStudy (anonymous):

The complete problem reads: Find the largest interval which includes x=0 for which the given initial-value problem has a unique solution. (Enter your answer in interval notation) (x-3)y"+2y=x, y(0)=0, y'(0)=1